how many years will it take $2600 to triple itself if it is invested at 3.6% simple interest
Answer by edjones(7569)

2600*.036x=3*2600
93.6x=7800
x=7800/93.6=83 1/3 yrs or 83 yrs 4 months
That is why you want compound interest.
Ed